OpenStudy (anonymous):

y varies jointly as the square x and the inverse of z. When y = 24 and x = 4, z = 6. Find the constant of variation and determine the value for x when y = 15 and z = 15. A.=135 B.=6.25 C.=√19 D.=5

OpenStudy (amistre64):

so y = kx and y = k/z together: y = kx/z, k = yz/x is constant

OpenStudy (amistre64):

might be good to define k in terms of x'y'z' as opposed to the other xyz part k = y'z'/x' is the initial stuff x = yz/k x = x' (yz)/(y'z') might be a better run